OpenStudy (anonymous):

Find the volume of the ellipsoid x^2+y^2+5z^2=100

OpenStudy (anonymous):

I think I remember using a double integral to find these . . .

OpenStudy (shubhamsrg):

is the ans 200 pi sqrt(20) ?

OpenStudy (anonymous):

no.

OpenStudy (unklerhaukus):

formula of an ellipsoid\[\left(\frac xa\right)^2+\left(\frac yb\right)^2+\left(\frac zc\right)^2=1\] where \(a,b,c ,\) are the semi-principal axes \[V_\text{{ellipsoid}}=\frac 43\pi abc\]

OpenStudy (shubhamsrg):

shouldnt it be since area of ellipse at at z=0 is pi(a.b) and total width along z axis is 2sqrt(20) should volume be area * width or 2pi.a.b.c @UnkleRhaukus where am i going wrong?